Grade 10Algebra

2x2+7x+1=1 find the number of integral values of x
….....

rosemead
7 Years agoGrade 10

1 Answer

Arun
7 Years ago
Dear student
 
it is solved as
 
2^(x^2 + 7x + 1) = 2^0
 
hence x^2 + 7x + 1 = 0
 
hence by sridharacharya formula
 
x = {– 7 ± sqrt(49 – 4) }/ 2
 
hence two solution exists